# Magnetic Field Response of Dipolar-Octupolar Quantum Spin Ice

### v1

@@ III.1 Generalities
$\mathcal{O}$ still maps the Ansatz to the same physical state if there exists a gauge transformation $U$ such that $U:\mathcal{H}_{\mathrm{GMFT}}(\{\bar{A}_{\mathcal{O}(\mathbf{r}_{\alpha}),\mathcal{O}(\mathbf{r}^{\prime}_{\beta})}\})\rightarrow\mathcal{H}_{\mathrm{GMFT}}(\{\bar{A}_{\mathbf{r}_{\alpha},\mathbf{r}^{\prime}_{\beta}}\})$. The PSG is then enlisted as a tool to list all inequivalent mean-field Ansätze that respect a given set of symmetries. In particular, it can enumerate all fully symmetric QSLs that respect all lattice symmetries. Notice that the PSG solution is dependent on the symmetries imposed. This implies that in the presence of a magnetic field, new fully symmetric states may be allowed if the field breaks some spatial symmetries.
@@ C.2.1 Algebraic Constraints
Under the [110] field, the Zeeman couplings on the four pyrochlore sites are $\sqrt{\frac{2}{3}}h(1,0,0,-1)$. The point group of the pyrochlore lattice can be considered the direct product of the sublattice permutation group $S_{4}$ and the transformation group generated by inversion (i.e., $O_{h}\simeq S_{4}\times\mathbb{Z}_{2}$). As shown in Appendix B, screw operator $S$ will map $S^{\pm}\rightarrow-S^{\pm}$. In doing so, it exchanges two pyrochlore sites. We can think of this as an odd permutation in the $S_{4}$ group. On the other hand, the generator $\bar{C}_{6}$ is an even permutation under $S_{4}$ since it is a three-cycle (i.e. (123)=(13)(12)). As such, all odd permutations will map $S^{\pm}\rightarrow-S^{\pm}$, whereas even permutations will not. This results in rather unexpected symmetry operations under finite magnetic fields. For example, under a [110] field, naively speaking, one would expect a mirror symmetry that swaps the 1st and the 2nd pyrochlore sites $\sigma_{12}=\bar{C}_{6}^{4}S\bar{C}_{6}(S\bar{C}_{6}^{-1})^{2}$, since both Zeeman couplings are 0. However, this operation is odd and maps $S^{\pm}\rightarrow-S^{\pm}$. As such, $\sigma_{12}:\sqrt{\frac{2}{3}}h(1,0,0,-1)\rightarrow\sqrt{\frac{2}{3}}h(-1,0,0,1)$ and therefore is not a symmetry.
@@ C.2.2 Solution to the PSG
$\phi_{I}(\mathbf{0}_{A})+\phi_{I}(\mathbf{0}_{B})-\psi_{IT_{3}}+2\phi_{\sigma}(\mathbf{0}_{A})-\psi_{\sigma T_{3}}=\psi_{\sigma I}$ (37e)
$\phi_{I}(\mathbf{0}_{A})+\phi_{I}(\mathbf{0}_{B})-\psi_{IT_{3}}+2\phi_{\sigma}(\mathbf{0}_{B})-2\psi_{\sigma T_{3}}=\psi_{\sigma I}.$ (37f)
By applying site-independent gauge transformation, we can fix $\psi_{\sigma T_{1}}=0$, which implies that $\psi_{IT_{3}}=\psi_{\sigma T_{3}}=\psi_{\sigma T_{2}}=0$. We can then make a sublattice-dependent gauge transformation such that $\phi_{\sigma}(\mathbf{0}_{A})=\psi_{\sigma}=\phi_{\sigma}(\mathbf{0}_{B})=0$, which implies that $\psi_{\sigma I}=\psi_{I}$. We can finally make a gauge transformation such that $\phi_{I}(\mathbf{0}_{A})=0$. Therefore, the final PSG solution is
@@ C.4.1 Algebraic Constraints
$\mathcal{G}_{T_{2}}^{-1}\left[T_{2}(\mathbf{r}_{\alpha})\right]\mathcal{G}_{\bar{C}_{4}}\left[T_{2}(\mathbf{r}_{\alpha})\right]\mathcal{G}_{T_{1}}^{-1}\left[\bar{C}_{4}^{-1}(\mathbf{r}_{\alpha})\right]\mathcal{G}_{\bar{C}_{4}}^{-1}$ $=e^{i\psi_{ST_{2}}}$ (46b)
$\mathcal{G}_{T_{3}}\mathcal{G}_{T_{2}}^{-1}\left[T_{2}T_{3}^{-1}(\mathbf{r}_{\alpha})\right]\mathcal{G}_{\bar{C}_{4}}\left[T_{2}T_{3}^{-1}(\mathbf{r}_{\alpha})\right]\mathcal{G}_{T_{2}}^{-1}\left[\bar{C}_{4}^{-1}(\mathbf{r}_{\alpha})\right]\mathcal{G}_{\bar{C}_{4}}^{-1}$ $=e^{i\psi_{ST_{3}}}$ (46c)
$\mathcal{G}_{T_{1}}\mathcal{G}_{T_{2}}^{-1}\left[T_{2}T_{1}^{-1}(\mathbf{r}_{\alpha})\right]\mathcal{G}_{\bar{C}_{4}}\left[T_{2}T_{1}^{-1}\right]\mathcal{G}_{T_{3}}^{-1}\left[\bar{C}_{4}^{-1}(\mathbf{r}_{\alpha})\right]\mathcal{G}_{\bar{C}_{4}}^{-1}$ $=e^{i\psi_{ST_{1}}}$ (46d)
$\mathcal{G}_{T_{2}}\mathcal{G}_{I}\left[T_{2}^{-1}(\mathbf{r}_{\alpha})\right]\mathcal{G}_{\bar{C}_{4}}\left[IT_{2}^{-1}(\mathbf{r}_{\alpha})\right]\mathcal{G}_{I}\left[I\bar{C}_{4}^{-1}(\mathbf{r}_{\alpha})\right]\mathcal{G}_{\bar{C}_{4}}^{-1}$ $=e^{i\psi_{SI}}$ (46e)
$\mathcal{G}_{T_{3}}^{-1}[T_{3}(\mathbf{r}_{\alpha})]\mathcal{G}_{S}[T_{3}(\mathbf{r}_{\alpha})]\mathcal{G}_{S}[S^{-1}T_{3}(\mathbf{r}_{\alpha})]\mathcal{G}_{I}[T_{3}S^{-2}(\mathbf{r}_{\alpha})]\mathcal{G}_{S}[S^{2}I(\mathbf{r}_{\alpha})]\mathcal{G}_{S}[SI(\mathbf{r}_{\alpha})]\mathcal{G}_{I}[I(\mathbf{r}_{\alpha})]$ $=e^{i\psi_{SIT}}.$ (46f)
@@ C.4.2 Solution to the PSG
$-\psi_{ST_{1}}r_{1}+\psi_{ST_{2}}r_{2}-\psi_{ST_{3}}r_{3}+n_{1}\pi(r_{1}+r_{3})\delta_{\alpha,B}+\psi_{\bar{C}_{4}}(\mathbf{0}_{\alpha}),$ (47)
@@
$2\psi_{IT_{1}}-\psi_{IT_{3}}+2\psi_{ST_{1}}$ $=0$ (49a)
$\psi_{IT_{1}}+\psi_{IT_{2}}-2\psi_{ST_{2}}$ $=0$ (49b)
$\psi_{IT_{1}}-\psi_{IT_{2}}+\psi_{IT_{3}}+2\psi_{ST_{3}}$ $=0$ (49c)
$\phi_{I}(\mathbf{0}_{A})+\phi_{I}(\mathbf{0}_{B})-\psi_{IT_{2}}+\psi_{ST_{2}}$ $=\psi_{SI}$ (49d)
$\psi_{IT_{1}}$ $=0$ (49e)
$\phi_{I}(\mathbf{0}_{A})+\phi_{I}(\mathbf{0}_{B})+4\phi_{S}(\mathbf{0}_{A})+\psi_{IT_{3}}+\psi_{ST_{1}}+2\psi_{ST_{2}}-2\psi_{ST_{3}}$ $=\psi_{SIT}$ (49f)
$\phi_{I}(\mathbf{0}_{A})+\phi_{I}(\mathbf{0}_{B})+4\phi_{S}(\mathbf{0}_{B})+2\psi_{IT_{3}}+\psi_{ST_{2}}-\psi_{ST_{3}}$ $=\psi_{SIT}$ (49g)
$\psi_{IT_{3}}-\psi_{ST_{1}}-\psi_{ST_{2}}+\psi_{ST_{3}}$ $=0.$ (49h)
The above relations give $\psi_{ST_{3}}=0$ and $\psi_{IT_{3}}=\psi_{IT_{2}}=\psi_{ST_{1}}=\psi_{ST_{2}}$. We can fix the gauge such that $\psi_{ST_{1}}=\psi_{SI}=\psi_{SIT}=0$. This implies that $\psi_{IT_{3}}=\psi_{IT_{2}}=\psi_{ST_{2}}=\psi_{ST_{3}}=0$. This further implies that $\psi_{I}=\psi_{SI}=0$. Finally, let us choose a gauge such that $\phi_{S}({\mathbf{0}_{A}})=\psi_{\bar{C}_{4}}=\phi_{S}(\mathbf{0}_{B})=0$ and $\phi_{I}(\mathbf{0}_{A})=\psi_{I}=\phi_{I}(\mathbf{0}_{B})=0$.
@@ C.7 PSG for Dipolar Spin Ice under Magnetic Fields
$\quad-\psi_{\sigma T_{2}}r_{2}+n_{S}\pi$ (55d)
$\phi_{I}(\mathbf{r}_{\alpha})$ $=\psi_{I}/2.$ (55e)
where $n_{S}\in\{0,1\}$, $\psi_{I}\in U(1)$. The PSG solutions under a $[001]$ field is:
@@
$\phi_{\bar{C}_{4}}(\mathbf{r}_{\alpha})=\frac{n_{1}\pi}{2}(-r_{1}(1+r_{1})+r_{3}(1+r_{3}))+n_{S}\pi(r_{2}+r_{3})$
$\qquad\quad+n_{1}\pi r_{1}r_{2}+(n_{1}\pi(r_{1}+r_{3}))\delta_{\alpha,B}+\psi_{\bar{C}_{4}}$ (56d)
$\phi_{I}(\mathbf{r}_{\alpha})=0.$ (56e)
where $n_{S}\in\mathbb{Z}_{2}$, $\psi_{\bar{C}_{4}}\in U(1)$. As such, the PSG solution for the dipolar case differs from that of the octupolar case. Despite this difference, both result in the same mean-field gauge configuration $\bar{A}_{\mathbf{r}_{\alpha}}$ as in Eqs. (52) and (54) up to a gauge transformation. For the XXZ model, this will not amount to a difference in the fluxes of hexagonal plaquettes (i.e., the MF Ansätze are equivalent). However, when considering the XYZ model, other mean-field parameters beyond the gauge field background are present. These may differ in the octupolar and dipolar spin ice cases, thus resulting in distinct MF Hamiltonians in both cases. In particular, there may be more states in the dipolar case as labeled by the additional PSG indices $n_{S}$ and $\psi_{\bar{C}_{4}}$.
@@ Appendix F Dynamical Spin Structure Factors
Figure 17: (a)-(b) Spinon dispersions and (c)-(d) dynamical spin structure factor in the global frame under a $[111]$ magnetic field for (a), (c) $J_{\pm}/J_{yy}=0.03$ and $h/J_{yy}=0.35$, and (b), (d) $J_{\pm}/J_{yy}=-0.03$, $h/J_{yy}=0.2$. The upper and lower edges of the two-spinon continuum are denoted by white lines.

### v2

@@ III.1 Generalities
$\mathcal{O}$ still maps the Ansätz to the same physical state if there exists a gauge transformation $U$ such that $U:\mathcal{H}_{\mathrm{GMFT}}(\{\bar{A}_{\mathcal{O}(\mathbf{r}_{\alpha}),\mathcal{O}(\mathbf{r}^{\prime}_{\beta})}\})\rightarrow\mathcal{H}_{\mathrm{GMFT}}(\{\bar{A}_{\mathbf{r}_{\alpha},\mathbf{r}^{\prime}_{\beta}}\})$. The PSG is then enlisted as a tool to list all inequivalent mean-field Ansätze that respect a given set of symmetries. In particular, it can enumerate all fully symmetric QSLs that respect all lattice symmetries. Notice that the PSG solution is dependent on the symmetries imposed. This implies that in the presence of a magnetic field, new fully symmetric states may be allowed if the field breaks some spatial symmetries.
@@ C.2.1 Algebraic Constraints
Under the [110] field, the Zeeman couplings on the four pyrochlore sites are $\sqrt{\frac{2}{3}}h(1,0,0,-1)$. The point group of the pyrochlore lattice can be considered the direct product of the sublattice permutation group $S_{4}$ and the transformation group generated by inversion (i.e., $O_{h}\simeq S_{4}\ltimes\mathbb{Z}_{2}$). As shown in Appendix B, screw operator $S$ will map $S^{\pm}\rightarrow-S^{\pm}$. In doing so, it exchanges two pyrochlore sites. We can think of this as an odd permutation in the $S_{4}$ group. On the other hand, the generator $\bar{C}_{6}$ is an even permutation under $S_{4}$ since it is a three-cycle (i.e. (123)=(13)(12)). As such, all odd permutations will map $S^{\pm}\rightarrow-S^{\pm}$, whereas even permutations will not. This results in rather unexpected symmetry operations under finite magnetic fields. For example, under a [110] field, naively speaking, one would expect a mirror symmetry that swaps the 1st and the 2nd pyrochlore sites $\sigma_{12}=\bar{C}_{6}^{4}S\bar{C}_{6}(S\bar{C}_{6}^{-1})^{2}$, since both Zeeman couplings are 0. However, this operation is odd and maps $S^{\pm}\rightarrow-S^{\pm}$. As such, $\sigma_{12}:\sqrt{\frac{2}{3}}h(1,0,0,-1)\rightarrow\sqrt{\frac{2}{3}}h(-1,0,0,1)$ and therefore is not a symmetry.
@@ C.2.2 Solution to the PSG
$\phi_{I}(\mathbf{0}_{A})+\phi_{I}(\mathbf{0}_{B})+\phi_{\sigma}(\mathbf{0}_{A})+\phi_{\sigma}(\mathbf{0}_{B})$
$\qquad\qquad\qquad\qquad\qquad-\psi_{IT_{3}}-\psi_{\sigma T_{3}}=\psi_{\sigma I}$ (37e)
By applying site-independent gauge transformation, we can fix $\psi_{\sigma T_{1}}=0$, which implies that $\psi_{IT_{3}}=\psi_{\sigma T_{3}}=\psi_{\sigma T_{2}}=0$ and $\phi_{\sigma}{\mathbf{0}_{A}}=\phi_{\sigma}(\mathbf{0}_{B})$. We can then make a sublattice-dependent gauge transformation such that $\phi_{\sigma}(\mathbf{0}_{A})=0=\psi_{\sigma}=\phi_{\sigma}(\mathbf{0}_{B})$, which implies that $\psi_{\sigma I}=\psi_{I}$. We can finally make a gauge transformation such that $\phi_{I}(\mathbf{0}_{A})=0$. Therefore, the final PSG solution is
@@ C.4.1 Algebraic Constraints
$\mathcal{G}_{T_{2}}^{-1}\left[T_{2}(\mathbf{r}_{\alpha})\right]\mathcal{G}_{\bar{C}_{4}}\left[T_{2}(\mathbf{r}_{\alpha})\right]\mathcal{G}_{T_{1}}^{-1}\left[\bar{C}_{4}^{-1}(\mathbf{r}_{\alpha})\right]\mathcal{G}_{\bar{C}_{4}}^{-1}$ $=e^{i\psi_{\bar{C}_{4}T_{2}}}$ (46b)
$\mathcal{G}_{T_{3}}\mathcal{G}_{T_{2}}^{-1}\left[T_{2}T_{3}^{-1}(\mathbf{r}_{\alpha})\right]\mathcal{G}_{\bar{C}_{4}}\left[T_{2}T_{3}^{-1}(\mathbf{r}_{\alpha})\right]\mathcal{G}_{T_{2}}^{-1}\left[\bar{C}_{4}^{-1}(\mathbf{r}_{\alpha})\right]\mathcal{G}_{\bar{C}_{4}}^{-1}$ $=e^{i\psi_{\bar{C}_{4}T_{3}}}$ (46c)
$\mathcal{G}_{T_{1}}\mathcal{G}_{T_{2}}^{-1}\left[T_{2}T_{1}^{-1}(\mathbf{r}_{\alpha})\right]\mathcal{G}_{\bar{C}_{4}}\left[T_{2}T_{1}^{-1}\right]\mathcal{G}_{T_{3}}^{-1}\left[\bar{C}_{4}^{-1}(\mathbf{r}_{\alpha})\right]\mathcal{G}_{\bar{C}_{4}}^{-1}$ $=e^{i\psi_{\bar{C}_{4}T_{1}}}$ (46d)
$\mathcal{G}_{T_{2}}\mathcal{G}_{I}\left[T_{2}^{-1}(\mathbf{r}_{\alpha})\right]\mathcal{G}_{\bar{C}_{4}}\left[IT_{2}^{-1}(\mathbf{r}_{\alpha})\right]\mathcal{G}_{I}\left[I\bar{C}_{4}^{-1}(\mathbf{r}_{\alpha})\right]\mathcal{G}_{\bar{C}_{4}}^{-1}$ $=e^{i\psi_{\bar{C}_{4}I}}$ (46e)
$\mathcal{G}_{T_{3}}^{-1}[T_{3}(\mathbf{r}_{\alpha})]\mathcal{G}_{\bar{C}_{4}}[T_{3}(\mathbf{r}_{\alpha})]\mathcal{G}_{\bar{C}_{4}}[\bar{C}_{4}^{-1}T_{3}(\mathbf{r}_{\alpha})]\mathcal{G}_{I}[T_{3}\bar{C}_{4}^{-2}(\mathbf{r}_{\alpha})]\mathcal{G}_{\bar{C}_{4}}[\bar{C}_{4}^{2}I(\mathbf{r}_{\alpha})]\mathcal{G}_{\bar{C}_{4}}[\bar{C}_{4}I(\mathbf{r}_{\alpha})]\mathcal{G}_{I}[I(\mathbf{r}_{\alpha})]$ $=e^{i\psi_{\bar{C}_{4}IT}}.$ (46f)
@@ C.4.2 Solution to the PSG
$+(\psi_{\bar{C}_{4}T_{2}}-\psi_{\bar{C}_{4}T_{1}}r_{1})+\psi_{\bar{C}_{4}T_{2}}r_{2}+(\psi_{\bar{C}_{4}T_{1}}-\psi_{\bar{C}_{4}T_{3}}r_{3})$
$+n_{1}\pi(r_{1}+r_{3})\delta_{\alpha,B}+\psi_{\bar{C}_{4}}(\mathbf{0}_{\alpha}),$ (47)
@@
$\psi_{IT_{1}}+\psi_{IT_{2}}-2\psi_{\bar{C}_{4}T_{2}}$ $=0$ (49a)
$2\psi_{IT_{1}}-\psi_{IT_{3}}+2\psi_{\bar{C}_{4}T_{1}}-2\psi_{\bar{C}_{4}T_{2}}$ $=0$ (49b)
$\psi_{IT_{1}}-\psi_{IT_{2}}+\psi_{IT_{3}}-2\psi_{\bar{C}_{4}T_{2}}+2\psi_{\bar{C}_{4}T_{3}}$ $=0$ (49c)
$\psi_{IT_{1}}-\psi_{IT_{2}}-\psi_{IT_{3}}+2\psi_{\bar{C}_{4}T_{1}}$ $=0$ (49d)
$-2\psi_{IT_{3}}+2\psi_{\bar{C}_{4}T_{1}}+2\psi_{\bar{C}_{4}T_{2}}-2\psi_{\bar{C}_{4}T_{3}}$ $=0$ (49e)
$\phi_{I}\left(\mathbf{0}_{A}\right)+\phi_{I}\left(\mathbf{0}_{B}\right)-\phi_{\bar{C}_{4}}\left(\mathbf{0}_{A}\right)+\phi_{\bar{C}_{4}}\left(\mathbf{0}_{B}\right)-\psi_{IT_{2}}+\psi_{\bar{C}_{4}T_{2}}$ $=\psi_{\bar{C}_{4}I}$ (49f)
$\phi_{I}\left(\mathbf{0}_{A}\right)+\phi_{I}\left(\mathbf{0}_{B}\right)+\phi_{\bar{C}_{4}}\left(\mathbf{0}_{A}\right)-\phi_{\bar{C}_{4}}\left(\mathbf{0}_{B}\right)-\psi_{IT_{1}}-\psi_{IT_{2}}+\psi_{\bar{C}_{4}T_{2}}$ $=\psi_{\bar{C}_{4}I}$ (49g)
$\phi_{I}\left(\mathbf{0}_{A}\right)+\phi_{I}\left(\mathbf{0}_{B}\right)+2\left(\phi_{\bar{C}_{4}}\left(\mathbf{0}_{A}\right)+\phi_{\bar{C}_{4}}\left(\mathbf{0}_{B}\right)\right)+\psi_{IT_{3}}+\psi_{\bar{C}_{4}T_{1}}+3\psi_{\bar{C}_{4}T_{2}}-2\psi_{\bar{C}_{4}T_{3}}$ $=\psi_{\bar{C}_{4}IT}$ (49h)
$\phi_{I}\left(\mathbf{0}_{A}\right)+\phi_{I}\left(\mathbf{0}_{B}\right)+2\left(\phi_{S}\left(\mathbf{0}_{A}\right)+\phi_{S}\left(\mathbf{0}_{B}\right)\right)+2\left(\psi_{IT_{3}}+\psi_{\bar{C}_{4}T_{2}}\right)-\psi_{\bar{C}_{4}T_{3}}$ $=\psi_{\bar{C}_{4}IT}$ (49i)
Let us fix the gauge such that $\phi_{\bar{C}_{4}}(\mathbf{0}_{A})=\phi_{\bar{C}_{4}}(\mathbf{0}_{B})$. This implies that $\psi_{IT_{1}}=0$. Then by fixing the gauge such that $\psi_{\bar{C}_{4}T_{1}}=\psi_{\bar{C}_{4}T_{3}}=0$, we enforce that $\psi_{\bar{C}_{4}T_{2}}=\psi_{IT_{1}}=\psi_{IT_{3}}=0$. Finally, by fixing $\psi_{\bar{C}_{4}I}=0$ and $\phi_{I}(\mathbf{0}_{A})=0$, we get that $\phi_{I}(\mathbf{0}_{A})=\phi_{I}(\mathbf{0}_{B})=0$.
@@ C.7 PSG for Dipolar Spin Ice under Magnetic Fields
$\quad-\psi_{\sigma T_{1}}r_{1}-\psi_{\sigma T_{2}}r_{2}$ (55d)
$\phi_{I}(\mathbf{r}_{\alpha})$ $=n_{I}\pi(r_{1}+r_{2})+\psi_{I}/2.$ (55e)
where $n_{I}\in\{0,1\}$. The PSG solutions under a $[001]$ field is:
@@
$\qquad\quad+n_{1}\pi r_{1}r_{2}+(n_{1}\pi(r_{1}+r_{3}))\delta_{\alpha,B}+\psi_{C}$ (56d)
$\phi_{I}(\mathbf{r}_{\alpha})=n_{I}\pi.$ (56e)
where $n_{I}\in\mathbb{Z}_{2}$, $\psi_{C}\in U(1)$. As such, the PSG solution for the dipolar case differs from that of the octupolar case. Despite this difference, both result in the same mean-field gauge configuration $\bar{A}_{\mathbf{r}_{\alpha}}$ as in Eqs. (52) and (54) up to a gauge transformation. For the XXZ model, this will not amount to a difference in the fluxes of hexagonal plaquettes (i.e., the MF Ansätze are equivalent). However, when considering the XYZ model, other mean-field parameters beyond the gauge field background are present. These may differ in the octupolar and dipolar spin ice cases, thus resulting in distinct MF Hamiltonians in both cases. In particular, there may be more states in the dipolar case as labeled by the additional PSG indices $n_{S}$ and $\psi_{\bar{C}_{4}}$.
@@ Appendix F Dynamical Spin Structure Factors
Figure 17: (a)-(b) Spinon dispersions and (c)-(d) dynamical spin structure factor in the global frame under a $[111]$ magnetic field for (a), (c) $J_{\pm}/J_{yy}=0.03$ and $h/J_{yy}=0.3$, and (b), (d) $J_{\pm}/J_{yy}=-0.03$, $h/J_{yy}=0.2$. The upper and lower edges of the two-spinon continuum are denoted by white lines.